# A Survey of Self-Evolving Agents: On Path to Artificial Super Intelligence

### v3

@@ 3.2.1 Memory Evolution
LLM-based agents are increasingly designed with long-term memory mechanisms that grow and adapt as the agent continues to solve tasks and interacts with its environment shan2025cognitive ; qian2023experiential . An evolving memory enables the agent to accumulate knowledge, recall past events, and adjust its behavior based on experience. Many works stress that effective memory management is crucial for agent performance zhong2024memorybank ; zhang2025g ; yan2024efficient . SAGE liang2024sage uses the Ebbinghaus forgetting curve to decide what to remember or forget. A-memxu2025mem updates the agent memory structure to create interconnected knowledge networks through dynamic indexing and linking, following the basic principles of the Zettelkasten method. Mem0 chhikara2025mem0 introduces a two-phase pipeline where the agent first extracts salient facts from recent dialogue and then decides how to update the long-term memory: the agent can ADD new facts, MERGE/UPDATE redundant ones, or DELETE contradictions. Such a mechanism ensures the agent’s long-term memory is coherent and up-to-date. MemInsight salama2025meminsight augments raw memories with semantic structure, which summarizes and tags past interactions for retrieval later. REMEMBER zhang2024large combines an LLM with a memory of experiences and uses reinforcement learning signals to decide how to update that memory after each episode.
@@
A critical aspect of memory evolution is enabling agents to learn heuristics or skills from past experiences. Rather than only retrieving exact past instances, advanced agents distill experiences into more general guidance zhao2024expel ; fu2024autoguide . Expel zhao2024expel processes past trajectories to generate insights and rules to guide further interactions. This experiential knowledge accumulation leads to measurable gains, as the agent steadily performs better with more experience. Other systems focus on storing higher-level building blocks of problem-solving. For instance, Agent Workflow Memory wang2024agent records common sub-task sequences (workflows) so that an agent solving a complex task can retrieve and reuse a proven sequence of actions rather than plan from scratch. In the Richelieu diplomacy agent, the system improves its negotiation strategies by augmenting its memory through self-play games, storing the insights from simulated interactions to refine future decisions guan2024richelieu . By generalizing from specific episodes to reusable knowledge, these approaches illustrate how memory evolution turns an agent’s one-time experiences into long-term competencies, which leads to agents evolving.
@@ 3.2.2 Prompt Optimization
While memory evolution focused on what knowledge an agent retains, Prompt Optimization (PO) enables LLM agents to self-evolve by refining the instructions it feeds to the backbone model, which directly alters the model’s behavior without modifying model weights ramnath2025systematic . Early research treats instruction design as a search problem. APE zhou2022large generates candidate prompts, scores them on validation examples, and selects the best. ORPO yang2023large extends this idea by letting the model iteratively rewrite its own prompt, guided by feedback on prior outputs. ADO lin2025ado introduces DSP that imposes semantic constraints on iteratively proposed prompts to facilitate finding the optimal prompt. ProTeGi pryzant2023automatic generates natural language "corrections" that are applied as edits to the prompt, forming a textual analogue of gradient descent. PromptAgent wang2023promptagent casts prompt discovery as Monte-Carlo Tree Search, exploring instruction space strategically, while evolutionary approaches like PromptBreeder fernando2023promptbreeder maintain a population to discover increasingly effective instructions. REVOLVE zhang2024revolve further stabilizes long optimization runs by tracking the trajectory of model responses and applying smoothed updates. Pushing this autonomy to its limit, SPO xiang2025self creates a fully self-contained loop where the model generates its training data and uses pairwise preference comparison on its outputs to refine the prompt, eliminating the need for any external labeled data or human feedback. Collectively, these techniques demonstrate that an agent can autonomously improve its prompting policy, turning prompt text into a learnable component that co-evolves with the agent’s experience.
@@
In complex systems, an agent often orchestrates a sequence of LLM calls or collaborates with other agents, making prompt design a multi-node problem. Frameworks such as DSPy represent an entire workflow as a graph whose sub-prompts are jointly tuned for a global objective khattab2023dspy . Trace wang2023trace , TextGrad yellamraju2024textgrad , and LLM-AutoDiff yin2025llm generalize this idea by treating each prompt as a parameter in a differentiable program and propagating natural-language “gradients” to refine every step. In collaborative scenarios, Multi-Agent System Search (MASS) zhou2025maasprompt first optimizes individual role prompts and then refines inter-agent communication patterns, while MAS-ZERO ke2025maszero dynamically proposes and revises role prompts to assemble an effective team for each new problem. Evolutionary systems such as EvoAgent yuan2024evoagent and AgentSquare shang2025agentsquare treat each agent along with prompts as the modules and use mutation and selection to discover specialized teams that outperform hand-crafted designs. These approaches extend PO from a single instruction to the language that defines whole workflows or societies of agents.
@@ 3.3 Tools
The primary impetus for autonomous tool creation is to overcome the inherent limitations of a fixed toolset, granting agents the flexibility to innovate on demand. Methodologies for this now span a spectrum from opportunistic discovery to formalized synthesis. At one end, agents like Voyager build an ever-expanding library of skills through emergent trial-and-error, driven by an intrinsic motivation to explore complex, open-ended environments like Minecraft wang2023voyager . This exploratory approach is powerful for generating a wide array of skills but may lack precision. In contrast, systems like Alita and ATLASS take a more reactive approach, often employing retrieval-augmented generation (RAG) to search open-source code repositories or write new functions from scratch the moment a capability gap is identified qiu2025alita ; Haque2025ATLASS . At the other end of the spectrum lie highly structured frameworks that treat tool creation as a deliberate engineering process. CREATOR, for example, disentangles abstract tool creation (e.g., reasoning about the general structure of a reusable function for averaging temperatures over N days) from concrete tool usage (e.g., deciding how to apply that function to a specific city and time range), which enhances modularity and reusability qian2023creator . Even more formally, SkillWeaver analyzes successful human or agent task trajectories to propose, synthesize, and hone new skills into robust, reusable APIs, ensuring a higher degree of initial quality zheng2025skillweaver . Furthermore, frameworks like CRAFT demonstrate that creating specialized toolsets for specific domains is essential to complement general-purpose models, enabling expert-level performance without sacrificing adaptability yuan2024craft . RL-GPT liu2024rl integrates generated code implementations into the RL pipeline, leveraging these as tools to tackle complex tasks while addressing simpler ones directly using a Code-as-Policy approach. This integration dynamically adapts and evolves in response to environmental feedback, enabling continuous improvement. However, this burgeoning autonomy introduces significant challenges, particularly around safety and security. The unconstrained generation of code risks creating tools with exploitable vulnerabilities or unintended harmful behaviors, making automated verification and sandboxing critical areas for future research.
@@
The proliferation of self-created tools necessitates a robust mechanism for their mastery; a newly generated tool is often a brittle script, not a reliable function. This is where iterative refinement becomes essential. Frameworks like LearnAct and From Exploration to Mastery establish a critical self-correction loop where the agent learns from its own experience zhao2024learnact ; fromexplorationtomastery . This involves tackling the difficult "credit assignment" problem: determining precisely which line of code or which parameter was responsible for a failure. To do this, the agent analyzes a rich variety of feedback signals—including compiler errors, unexpected API return values, environmental state changes, or even implicit signals from a user’s subsequent actions. The goal is not only to debug the tool’s underlying code but also to refine its documentation (e.g., its docstring and argument descriptions), which is crucial for improving the agent’s ability to understand and correctly use the tool in the future. This refinement process also opens the door for valuable human-agent collaboration. While full autonomy is the ultimate goal, many systems can be designed with a "human in the loop," where a human expert can provide corrections, offer high-level suggestions, or validate a newly created tool. This collaborative approach can significantly accelerate the mastery process and ensure that the agent’s skills align with human intentions and safety standards. Ultimately, this self-honing process is what elevates a nascent skill into a dependable capability, ensuring the agent’s growing skill library increases not just in quantity, but more importantly, in quality and robustness.
@@
As an agent’s mastered skill library grows into the hundreds or thousands, it faces a "curse of abundance." The challenge shifts from creating tools to efficiently managing and selecting from them. A large library creates a massive search space, making traditional retrieval methods slow and inaccurate. To overcome this, ToolGen represents a fundamental paradigm shift by encoding tools as unique tokens within the language model’s vocabulary. This elegantly reframes tool retrieval as a generation problem, leveraging the transformer’s immense pattern-recognition capabilities to predict the most appropriate tool as a natural continuation of its thought process wang2025toolgen . Beyond selecting a single tool, advanced agents must also excel at tool composition—learning to chain multiple tools in novel sequences to solve multi-step problems. This is a higher-order management task. Architectural approaches like AgentSquare engage in a form of meta-learning, automatically searching the modular design space of an agent—including its planning, memory, and tool-use components—to find an optimal configuration for complex task execution shang2025agentsquare . As a logical endpoint to this evolutionary trend, visionary concepts like the Darwin Godel Machine propose a framework for open-ended evolution, where the agent can fundamentally rewrite its own core code. In this vision, the distinction between the agent and its tools blurs, leading to a recursive cascade of self-improvement that transcends tool enhancement alone zhang2025darwin . In essence, this entire evolutionary path aims to establish a closed and virtuous cycle: a truly autonomous agent that can perceive gaps in its capabilities, create novel solutions, master them through practice, and seamlessly integrate them into a coherently managed and ever-expanding repertoire.
@@ 3.4 Architecture
The defining feature of next-generation agentic systems is their intrinsic capacity for self-improvement. This marks a fundamental shift from systems with fixed capabilities to those that can autonomously enhance their performanceliu2025advances . By treating their own internal logic and collaborative structures as optimizable components, these systems can adapt their behavior and design in response to feedback, achieving a level of efficiency and effectiveness that static designs cannot match. This section details how this self-optimization is realized, first by examining improvements within single-agent systems and then by exploring the co-evolution of complex multi-agent systems.
@@ 3.4.1 Single-Agent System Optimization
Optimizing a single LLM call is straightforward in isolation, but within an agentic system, it becomes a difficult credit assignment problem, as the effect of any single change is obscured by subsequent steps. Research addresses this by making node-level components optimizable, following two main strategies. The first focuses on refining nodes within a fixed agentic topology. A prime example is TextGrad yellamraju2024textgrad , which, inspired by backpropagation, uses "textual gradients" to propagate feedback from the final output backward through the workflow, guiding systematic, local refinements at each node without altering the system’s overall structure. The second, parallel strategy integrates this component-level optimization directly into the search for the system’s architecture itself. Under this approach, node characteristics become tunable parameters in a larger search space. For instance, frameworks can embed prompt engineering directly into the search loop, allowing the system to discover not just the optimal workflow but also the most effective instruction for each agent simultaneously zhou2025maasprompt . Similarly, EvoFlow zhang2025evoflow uses evolutionary algorithms to construct heterogeneous workflows by selecting the most suitable LLM for each task from a diverse pool. This holistic strategy enables the discovery of systems that are co-optimized for both their structure and individual agent capabilities, effectively balancing metrics like overall performance and cost ye2025xmas .
@@
Building upon the optimization of individual LLM-invoking nodes, a more profound level of self-improvement targets the autonomous agent as a holistic entity. This evolution proceeds along two main fronts: optimizing the agent’s high-level architectural design and enabling the agent to directly modify its own source code. The first approach focuses on discovering the optimal agent structure. AgentSquare shang2025agentsquare exemplifies this by defining a modular design space of components like planners and memory modules, then using an evolutionary algorithm to find the most effective combination for a given task. The second front involves agents that dynamically rewrite their own operational code. This is seen in radical systems like the Darwin Gödel Machine zhang2025darwin , which recursively modifies its own Python codebase, and AlphaEvolve novikov2025alphaevolve , which uses evolutionary coding to improve specific algorithms. Similarly, Gödel Agent yin2025godelagentselfreferentialagent provides a self-referential framework for agents to analyze and alter their logic. Together, these two directions (optimizing the agent’s architectural “blueprint” and its functional code) demonstrate a key trend toward turning the agent’s fundamental structure and logic into learnable components.

### v4

@@ 3.2.1 Memory Evolution
LLM-based agents are increasingly designed with long-term memory mechanisms that grow and adapt as the agent continues to solve tasks and interacts with its environment(Shan et al., 2025; Qian et al., 2023a). An evolving memory enables the agent to accumulate knowledge, recall past events, and adjust its behavior based on experience. Many works stress that effective memory management is crucial for agent performance(Zhong et al., 2024; Zhang et al., 2025e; Yan et al., 2024). SAGE(Liang et al., 2024) uses the Ebbinghaus forgetting curve to decide what to remember or forget. A-mem(Xu et al., 2025a) updates the agent memory structure to create interconnected knowledge networks through dynamic indexing and linking, following the basic principles of the Zettelkasten method. Mem0(Chhikara et al., 2025) introduces a two-phase pipeline where the agent first extracts salient facts from recent dialogue and then decides how to update the long-term memory: the agent can ADD new facts, MERGE/UPDATE redundant ones, or DELETE contradictions. Furthermore, Memory-R1 (Yan et al., 2025) presents a reinforcement learning framework to train a dedicated Memory Manager agent that learns to select structured operations like ADD, UPDATE, and DELETE. Such a mechanism ensures the agent’s long-term memory is coherent and up-to-date. MemInsight(Salama et al., 2025) augments raw memories with semantic structure, which summarizes and tags past interactions for retrieval later. REMEMBER(Zhang et al., 2024a) combines an LLM with a memory of experiences and uses reinforcement learning signals to decide how to update that memory after each episode. Memento (Zhou et al., 2025b) enables continual adaptation without fine-tuning the LLM’s parameters by employing online reinforcement learning to optimize a case-retrieval policy, which allows the agent to learn from past experiences stored in an evolving memory bank. MemGen (Zhang et al., 2025f) introduces a dynamic generative memory that operates in a latent space. It uses a learned memory trigger to decide when to invoke memory and a weaver to construct latent token sequences, enabling a fluid interweaving of reasoning and memory.
@@
A critical aspect of memory evolution is enabling agents to learn heuristics or skills from past experiences. Rather than only retrieving exact past instances, advanced agents distill experiences into more general guidance(Zhao et al., 2024a; Fu et al., 2024). Expel(Zhao et al., 2024a) processes past trajectories to generate insights and rules to guide further interactions. This experiential knowledge accumulation leads to measurable gains, as the agent steadily performs better with more experience. ReasoningBank (Ouyang et al., 2025) further develops this idea by distilling generalizable reasoning strategies from both successful and failed experiences into a structured memory. It also introduces memory-aware test-time scaling to generate diverse experiences on each task. Other systems focus on storing higher-level building blocks of problem-solving. For instance, Agent Workflow Memory(Wang et al., 2024j) records common sub-task sequences (workflows) so that an agent solving a complex task can retrieve and reuse a proven sequence of actions rather than plan from scratch. Similarly, MUSE (Yang et al., 2025a) introduces an experience-driven agent for long-horizon tasks, centered on a hierarchical memory module that organizes experience into strategic, procedural, and tool-use memories. The agent populates this memory through a Plan-Execute-Reflect-Memorize loop, enabling it to learn on the job. In the Richelieu diplomacy agent, the system improves its negotiation strategies by augmenting its memory through self-play games, storing the insights from simulated interactions to refine future decisions(Guan et al., 2024). By generalizing from specific episodes to reusable knowledge, these approaches illustrate how memory evolution turns an agent’s one-time experiences into long-term competencies, which leads to agents evolving.
@@ 3.2.2 Prompt Optimization
While memory evolution focused on what knowledge an agent retains, Prompt Optimization (PO) enables LLM agents to self-evolve by refining the instructions it feeds to the backbone model, which directly alters the model’s behavior without modifying model weights(Ramnath et al., 2025). Early research treats instruction design as a search problem. APE(Zhou et al., 2022) generates candidate prompts, scores them on validation examples, and selects the best. ORPO(Yang et al., 2023) extends this idea by letting the model iteratively rewrite its own prompt, guided by feedback on prior outputs. ADO(Lin et al., 2025a) introduces DSP that imposes semantic constraints on iteratively proposed prompts to facilitate finding the optimal prompt. ProTeGi(Pryzant et al., 2023) generates natural language "corrections" that are applied as edits to the prompt, forming a textual analogue of gradient descent. PromptAgent(Wang et al., 2023d) casts prompt discovery as Monte-Carlo Tree Search, exploring instruction space strategically, while evolutionary approaches like PromptBreeder(Fernando et al., 2023) maintain a population to discover increasingly effective instructions. REVOLVE(Zhang et al., 2024e) further stabilizes long optimization runs by tracking the trajectory of model responses and applying smoothed updates. Pushing this autonomy to its limit, SPO(Xiang et al., 2025) creates a fully self-contained loop where the model generates its training data and uses pairwise preference comparison on its outputs to refine the prompt, eliminating the need for any external labeled data or human feedback. Collectively, these techniques demonstrate that an agent can autonomously improve its prompting policy, turning prompt text into a learnable component that co-evolves with the agent’s experience. To address the brevity bias and context collapse of some optimizers, Agentic Context Engineering (ACE) (Zhang et al., 2025m) treats contexts as comprehensive playbooks that accumulate strategies over time. It uses a modular agentic process with incremental updates to evolve contexts for both offline prompt optimization and online memory adaptation.
@@
In complex systems, an agent often orchestrates a sequence of LLM calls or collaborates with other agents, making prompt design a multi-node problem. Frameworks such as DSPy represent an entire workflow as a graph whose sub-prompts are jointly tuned for a global objective(Khattab et al., 2023). Trace(Wang et al., 2023c), TextGrad(Yellamraju and others, 2024), and LLM-AutoDiff(Yin and Wang, 2025) generalize this idea by treating each prompt as a parameter in a differentiable program and propagating natural-language “gradients” to refine every step. In collaborative scenarios, Multi-Agent System Search (MASS)(Zhou et al., 2025a) first optimizes individual role prompts and then refines inter-agent communication patterns, while MAS-ZERO(Ke et al., 2025) dynamically proposes and revises role prompts to assemble an effective team for each new problem. Evolutionary systems such as EvoAgent(Yuan et al., 2024b) and AgentSquare(Shang et al., 2025) treat each agent along with prompts as the modules and use mutation and selection to discover specialized teams that outperform hand-crafted designs. These approaches extend PO from a single instruction to the language that defines whole workflows or societies of agents.
@@ 3.3 Tools
The primary impetus for autonomous tool creation is to overcome the inherent limitations of a fixed toolset, granting agents the flexibility to innovate on demand. Methodologies for this now span a spectrum from opportunistic discovery to formalized synthesis. At one end, agents like Voyager build an ever-expanding library of skills through emergent trial-and-error, driven by an intrinsic motivation to explore complex, open-ended environments like Minecraft(Wang et al., 2023a). This exploratory approach is powerful for generating a wide array of skills but may lack precision. In contrast, systems like ATLASS, Alita, and Live-SWE-Agent take a more reactive approach, often creating new tools from scratch or employing retrieval-augmented generation (RAG) to search open-source code repositories the moment a capability gap is identified(Haque et al., 2025; Qiu et al., 2025b; a; Xia et al., 2025a). At the other end of the spectrum lie highly structured frameworks that treat tool creation as a deliberate engineering process. CREATOR, for example, disentangles abstract tool creation (e.g., reasoning about the general structure of a reusable function for averaging temperatures over N days) from concrete tool usage (e.g., deciding how to apply that function to a specific city and time range), which enhances modularity and reusability(Qian et al., 2023b). Even more formally, SkillWeaver analyzes successful human or agent task trajectories to propose, synthesize, and hone new skills into robust, reusable APIs, ensuring a higher degree of initial quality(Zheng et al., 2025a). Furthermore, frameworks like CRAFT demonstrate that creating specialized toolsets for specific domains is essential to complement general-purpose models, enabling expert-level performance without sacrificing adaptability(Yuan et al., 2024a). RL-GPT(Liu et al., 2024) integrates generated code implementations into the RL pipeline, leveraging these as tools to tackle complex tasks while addressing simpler ones directly using a Code-as-Policy approach. This integration dynamically adapts and evolves in response to environmental feedback, enabling continuous improvement. However, this burgeoning autonomy introduces significant challenges, particularly around safety and security. The unconstrained generation of code risks creating tools with exploitable vulnerabilities or unintended harmful behaviors, making automated verification and sandboxing critical areas for future research.
@@
The proliferation of self-created tools necessitates a robust mechanism for their mastery; a newly generated tool is often a brittle script, not a reliable function. This is where iterative refinement becomes essential. Frameworks like LearnAct and From Exploration to Mastery establish a critical self-correction loop where the agent learns from its own experience(Zhao et al., 2024b; Qu et al., 2025). This involves tackling the difficult "credit assignment" problem: determining precisely which line of code or which parameter was responsible for a failure. To do this, the agent analyzes a rich variety of feedback signals—including compiler errors, unexpected API return values, environmental state changes, or even implicit signals from a user’s subsequent actions. The goal is not only to debug the tool’s underlying code but also to refine its documentation (e.g., its docstring and argument descriptions), which is crucial for improving the agent’s ability to understand and correctly use the tool in the future. This refinement process also opens the door for valuable human-agent collaboration. While full autonomy is the ultimate goal, many systems can be designed with a "human in the loop," where a human expert can provide corrections, offer high-level suggestions, or validate a newly created tool. This collaborative approach can significantly accelerate the mastery process and ensure that the agent’s skills align with human intentions and safety standards. Ultimately, this self-honing process is what elevates a nascent skill into a dependable capability, ensuring the agent’s growing skill library increases not just in quantity, but more importantly, in quality and robustness.
@@
As an agent’s mastered skill library grows into the hundreds or thousands, it faces a "curse of abundance." The challenge shifts from creating tools to efficiently managing and selecting from them. A large library creates a massive search space, making traditional retrieval methods slow and inaccurate. To overcome this, ToolGen (Wang et al., 2025j) represents a fundamental paradigm shift by encoding tools as unique tokens within the language model’s vocabulary. This elegantly reframes tool retrieval as a generation problem, leveraging the transformer’s immense pattern-recognition capabilities to predict the most appropriate tool as a natural continuation of its thought process. TOOLMEM (Xiao et al., 2025b) enables agents to learn and store the strengths and weaknesses of different tools in a dedicated memory. At inference, the agent retrieves this knowledge to make more informed decisions, optimizing tool selection for specific task requirements. Beyond selecting a single tool, advanced agents must also excel at tool composition—learning to chain multiple tools in novel sequences to solve multi-step problems. This is a higher-order management task. Architectural approaches like AgentSquare engage in a form of meta-learning, automatically searching the modular design space of an agent—including its planning, memory, and tool-use components—to find an optimal configuration for complex task execution(Shang et al., 2025). As a logical endpoint to this evolutionary trend, visionary concepts like the Darwin Godel Machine propose a framework for open-ended evolution, where the agent can fundamentally rewrite its own core code. In this vision, the distinction between the agent and its tools blurs, leading to a recursive cascade of self-improvement that transcends tool enhancement alone(Zhang et al., 2025h). In essence, this entire evolutionary path aims to establish a closed and virtuous cycle: a truly autonomous agent that can perceive gaps in its capabilities, create novel solutions, master them through practice, and seamlessly integrate them into a coherently managed and ever-expanding repertoire.
@@ 3.4 Architecture
The defining feature of next-generation agentic systems is their intrinsic capacity for self-improvement. This marks a fundamental shift from systems with fixed capabilities to those that can autonomously enhance their performance(Liu et al., 2025b). By treating their own internal logic and collaborative structures as optimizable components, these systems can adapt their behavior and design in response to feedback, achieving a level of efficiency and effectiveness that static designs cannot match. This section details how this self-optimization is realized, first by examining improvements within single-agent systems and then by exploring the co-evolution of complex multi-agent systems.
@@ 3.4.1 Single-Agent System Optimization
Optimizing a single LLM call is straightforward in isolation, but within an agentic system, it becomes a difficult credit assignment problem, as the effect of any single change is obscured by subsequent steps. Research addresses this by making node-level components optimizable, following two main strategies. The first focuses on refining nodes within a fixed agentic topology. A prime example is TextGrad (Yellamraju and others, 2024), which, inspired by backpropagation, uses "textual gradients" to propagate feedback from the final output backward through the workflow, guiding systematic, local refinements at each node without altering the system’s overall structure. The second, parallel strategy integrates this component-level optimization directly into the search for the system’s architecture itself. Under this approach, node characteristics become tunable parameters in a larger search space. For instance, frameworks can embed prompt engineering directly into the search loop, allowing the system to discover not just the optimal workflow but also the most effective instruction for each agent simultaneously (Zhou et al., 2025a). Similarly, EvoFlow (Zhang et al., 2025c) uses evolutionary algorithms to construct heterogeneous workflows by selecting the most suitable LLM for each task from a diverse pool. This holistic strategy enables the discovery of systems that are co-optimized for both their structure and individual agent capabilities, effectively balancing metrics like overall performance and cost (Ye et al., 2025).
@@
Building upon the optimization of individual LLM-invoking nodes, a more profound level of self-improvement targets the autonomous agent as a holistic entity. This evolution proceeds along two main fronts: optimizing the agent’s high-level architectural design and enabling the agent to directly modify its own source code. The first approach focuses on discovering the optimal agent structure. AgentSquare (Shang et al., 2025) exemplifies this by defining a modular design space of components like planners and memory modules, then using an evolutionary algorithm to find the most effective combination for a given task. The second front involves agents that dynamically rewrite their own operational code. This is seen in radical systems like the Darwin Gödel Machine (Zhang et al., 2025h), which recursively modifies its own Python codebase, and AlphaEvolve (Novikov et al., 2025), which uses evolutionary coding to improve specific algorithms. Similarly, Gödel Agent (Yin et al., 2025) provides a self-referential framework for agents to analyze and alter their logic. Together, these two directions (optimizing the agent’s architectural “blueprint” and its functional code) demonstrate a key trend toward turning the agent’s fundamental structure and logic into learnable components. MemEvolve (Zhang et al., 2025g) introduces a meta-evolutionary framework that evolves not just the agent’s experiential memory, but the memory system’s architecture itself. Through a bilevel optimization process, it adapts the mechanisms for encoding, storing, and retrieving information to better suit specific task domains.